Suppose that the rate of interest is 6.25%. The following annual net returns are expected on an investment project in the next five years.: $10,500, $11,200, $12,400, $8,640 and $5,530, respectively. The cost of the project is $15,300. Calculate NPV.
Answer #1

ANSWER:

I = 6.25%

n = 5 years

initial cost = 15,300

npv = initial cost + annual return in 1st year(p/f,i,n) + annual return in 1st year(p/f,i,n) + annual return in 1st year(p/f,i,n) + annual return in 1st year(p/f,i,n) + annual return in 1st year(p/f,i,n)

npv = -15,300 + 10,500(p/f,6.25%,1) + 11,200(p/f,6.25%,2) + 12,400(p/f,6.25%,3) + 8,640(p/f,6.25%,4) + 5,530(p/f,6.25%,5)

npv = -15,300 + 10,500 * 0.9412 + 11,200 * 0.8858 + 12,400 * 0.8337 + 8,640 * 0.7847 + 5,530 * 0.7385

npv = -15,300 + 9,882.35 + 9,921.11 + 10,337.96 + 6,779.51 + 4,083.95

npv = 25,704.88

so the net present value of the project is $25,704.88